# HSIDMamba: Exploring Bidirectional State-Space Models for Hyperspectral Denoising

### v1

@@ III-B2 Hyperspectral Continuous Scan Block
where $F_{j}$ represents the reordered 2D features, and t represents the number of rearranged sequences.
@@ III-C Complexity Analysis
In previous work, CNNs and Transformers were two common architectures that played important roles in processing sequence and image data. Compared to traditional CNNs and Transformers, HSDM takes advantage of linear complexity. The time complexity of traditional CNNs depends on the size of the input image and the number of filters, while the time complexity of Transformers is quadratic with the length of the input sequence. Here we analyze their time complexity.
@@
where $T$ represents the number of times to rearrange 2D features.
@@ IV Experiments
In this section, we explore both synthetic, real, and remote datasets being corrupted by mixed noise. We extensively assess our proposed HSDM and conduct experiments to analyze its efficacy thoroughly.
IV-A Overall Network Architecture
As illustrated in Figure 3, in the baseline configuration of HSDT, we employed a total of 3 layers of HCSL, each comprising 4 Blocks for alternating forward and backward Continuous Scan. Additionally, prior to the modules, a residual connection and feature concatenation from the module output were utilized. Finally, skip connections were employed to add the initial input of each layer to the overall layer output, ensuring the stability of the denoising pattern.
IV-A1 Training Datasets
In training our proposed HSDM, we adopt a dataset configuration similar to that of [30], where 100 images are randomly sampled from the ICVL dataset [31]. These images are centrally cropped to achieve a spatial resolution of 512×512 with 31 spectral bands. Further adjustments are made to the spatial resolution of the images to 64×64 through operations including random cropping, rotation, and flipping, resulting in approximately 53,000 training samples.
IV-A2 Testing Datasets
To evaluate the model’s generalization, testing datasets also encompass remote sensing datasets including the Washington DC dataset, real datasets: Urban [33].
IV-A3 Comparison Methods and Evaluation Metrics
Our experiments encompass a comparative analysis between HSDM and a spectrum of methodologies, including traditional model-based approaches including LRTDTV [34] and LLRGTV [7], as well as deep learning-based methods like GRNet [35], MACNET [24], T3SC [36], SST [37], SERT [27], and HSDT [38]. Notably, we specifically focus on contrasting our architecture against SST, SERT, and HSDT, which are all Transformer-based methods, aiming to underscore the denoising advantages inherent in our approach. To comprehensively evaluate the performance of these methodologies, we employ three distinct image quality evaluation metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM) [39], and Spectral Angle Mapper (SAM) [40]. Through this meticulous comparison, we aim to elucidate the superior denoising capabilities of HSDM in various real-world scenarios.
@@ IV-A3 Comparison Methods and Evaluation Metrics
HSDM 46.31 0.9982 0.0289 42.74 0.9964 0.0362 40.77 0.9943 0.0408 39.39 0.9922 0.0447
Noisy LRTDTV LLRGTV MAC-Net T3SC
SST SERT HSDT HSDM GT
@@
HSDM (Ours) 43.44 0.9967 0.0371 43.03 0.9964 0.0388 42.99 0.9964 0.0391 42.43 0.9958 0.0418 41.30 0.9950 0.0459
@@ IV-A5 Implementation Details
We introduce the random addition of complex noise from the first four noises during network training to improve the network’s adaptability and resilience. HSDM utilizes the Adam optimizer to minimize the mean squared error, following a training strategy similar to that outlined in [30], with minor modifications in the learning rate schedule. Specifically, the learning rate undergoes three stages: initially set to 3e-4, it is then reduced by a factor of 5 after every 20 epochs. Training is performed with a batch size of 8 on an RTX3090 for 100 epochs.
@@ IV-B Efficiency analysis
In this subsection, we analyze the performance of different methods regarding model parameter size and time consumption cost on the CAVE dataset with noise case 5. All test experiments are conducted on a 3090 GPU, and the results are presented in Table III.
TABLE III: Model complexity comparisons on the CAVE dataset with dimensions of 512×512×31.
@@
Noisy GRNet QRNN3D T3SC GRUNet
SST SERT HSDT HSDM GT
@@ IV-C Experimental Results
To ensure the effectiveness of the comparative methods, for model-based approaches, we select corresponding test settings, and for deep learning-based methods, we compare them with their publicly available pre-trained models. All models are trained on the ICVL dataset. For datasets with a larger number of spectral bands, some comparative methods are only applicable to data with 31 spectral bands. We employ a sliding window strategy for denoising, and the average of the results within all windows is used as the final result.
IV-C1 Complex Noise on Synthetic Datasets
In this section, we investigate the denoising performance of various methods under fixed Gaussian noise with different standard deviations ($\sigma$) ranging from 10 to 70, applied to the original HSI dataset. Our method is meticulously evaluated against baseline approaches on the ICVL dataset, with the comparative outcomes meticulously detailed in Table I. Furthermore, to provide a comprehensive visual assessment, Figure 4 displays the denoising results obtained using different methods, among which HSDM exhibits superior performance overall, particularly evident at the noise level of $\sigma=70$.
Upon scrutiny of the visual results, it is discerned that while LRTDTV, LLRGTV, and MAC-Net exhibit discernible pseudo-artifacts under the imposed Gaussian noise, our proposed HSDM method demonstrates superior efficacy in effectively eliminating Gaussian noise artifacts. Notably, HSDM adeptly preserves intricate details within the HSI dataset, thus ensuring superior fidelity and accuracy compared to alternative denoising techniques.
@@ IV-C2 Complex Noise on Synthetic Datasets
Furthermore, visual comparison results under mixed noise settings on the ICVL dataset are presented in Figure 5. It is noteworthy to observe a spectral shift phenomenon in denoising outcomes obtained with GRNet, wherein certain pixels transition from yellow to blue hues. Similarly, T3SC exhibits noticeable color distortion issues. In stark contrast, our proposed MambaHSID model maintains superior spectral consistency and visual fidelity, closely aligning with the GroundTruth across both spectral and spatial domains.
@@
Noisy LRTDTV LLRGTV MAC-Net T3SC
@@
QRNN3D SST SERT HSDT HSDM
Figure 6: Visual comparison of denoising on the WDC dataset under mixture noise. Bands 76, 43, and 10 combine to create the visual image.
@@
Noisy LRTDTV LLRGTV MAC-Net T3SC
@@
QRNN3D SST SERT HSDT HSDM
Figure 7: Visual comparison of denoising on the Urban dataset. The HSI bands 102,138 and 202 combine to create the visual pseudocolor image.
@@ IV-C3 Noise on Real World
To validate the model’s robustness under real-world noise conditions, we conducted hyperspectral image denoising experiments on representative datasets Urban. Figures 7 illustrate the visual denoising performance across Urban datasets.
Moreover, we extended our denoising experiments to include the CAVE dataset, and the quantitative outcomes are detailed in Table IV. HSDM showcased remarkable superiority in real noise removal, as evidenced by its higher PSNR and SSIM values compared to other methods. While LRTDTV and LLRGTV demonstrated a certain degree of alignment with the SAM metric changes on real datasets due to their priori global low-rank assumptions, their performance was relatively poorer on additional datasets. In contrast, HSDM consistently exhibited superior performance across a wider range of datasets, underscoring its versatility and adaptability to various HSI denoising scenarios.
@@ IV-D1 Effectiveness of scanning mechanism
To better explore the impact of scanning mechanisms combined with SSM on HSI denoising, we conducted experiments using both popular Sweep scanning and adjacent scanning methods on different datasets. We performed both unidirectional and bidirectional experiments for each scanning method, and the results are summarized in Table IV. The findings provide evidence supporting the effectiveness of our bidirectional adjacent scanning approach.
@@
Baseline Params(M) ICVL CAVE WDC
@@
Results demonstrate that our bidirectional adjacent scanning approach outperforms other scanning methods, indicating its effectiveness in capturing complex spatial-spectral relationships within the hyperspectral data. This finding underscores the importance of considering bidirectional scanning mechanisms when implementing state space models for HSI denoising.
@@ IV-D2 Effectiveness of modifications
IV-D3 Comparison of the hidden dims
The impact of varying numbers of hidden channels on the model’s feature-capturing ability is significant. We analyzed different numbers of hidden channels, presented in Table VI. This decision strikes a better balance between the model’s efficiency and performance considerations. By increasing the number of hidden channels to 48, the model gains enhanced capacity to capture and represent complex features within the data. This decision contributes to improved model performance without excessively increasing computational complexity, thus optimizing the trade-off between efficiency and effectiveness.
@@ IV-D3 Comparison of the hidden dims
hidden dim Params(M) PSNR(dB) SSIM SAM
24 0.24 38.85 0.9901 0.0418
36 0.44 39.71 0.9936 0.0350
48 0.68 41.30 0.9950 0.0459
64 1.08 41.50 0.9937 0.0374
96 2.17 41.76 0.9946 0.0360
@@ IV-E Parameter analysis
HSDM consistently maintains its GPU memory consumption and FPS rate, even with gradually increasing blocks. In contrast, the primary baseline Bidirectional HSDM experiences a decrease in FPS due to increased parameters. Notably, HSID Mamba, with blocks of 4, performs faster without altering parameters. Study on block sizes ranging from 2, 4, 6, to 8 reveals a deterioration, highlighting the importance of block sizes for optimal performance. Too few blocks may reduce effective feature extraction, while too many may lead to feature redundancy. Targeted selection of the block size balances computational efficiency and model effectiveness, ensuring effective feature capture while maintaining efficient processing speeds.
@@ V Conclusion
In this paper, we present HSIDMamba(HSDM), an innovative HSI denoising approach grounded in the selective state space model framework. HSDM is engineered to tackle the unique obstacles associated with processing long sequences inherent in HSI. By seamlessly integrating convolution and attention mechanisms, HSDM excels at capturing intricate spatial-spectral dependencies with remarkable computational efficiency, thanks to its linear complexity. Moreover, it introduces a novel bidirectional continuous scanning mechanism meticulously tailored to the nuances of HSI, thereby amplifying its ability to perceive and process spatial-spectral information effectively. Through extensive experimental validations conducted across diverse datasets, HSDM surpasses existing methodologies, highlighting its effectiveness and efficiency in tackling the challenges of HSI denoising.

### v2

@@ III-B3 Hyperspectral Continuous Scan Block
where $F_{j}$ represents the reordered 2D features, and t represents the number of rearranged sequences. When $i$ is odd, it indicates forward scanning, when $i$ is even, it signifies backward.
@@ III-C Complexity Analysis
In previous work, CNNs and Transformers played important roles in processing sequence and image data. However, their time complexities present optimization challenges. For CNNs, complexity grows with input image size and filter count—optimizations can include using separable convolutions to reduce parameters. Transformers exhibit quadratic complexity relative to input length but can be optimized through attention mechanisms like sparse attention. Our proposed HSDM, on the contrary, benefits from linear complexity, mitigating scalability issues inherent to the other architectures. This subsection aims to analyze the time complexity across these base architectures.
@@
where ${\rm{T}}(\ll C)$ represents the number of times to rearrange 2D features.
@@ IV Experiments
HSDM-B 46.31 0.9982 0.0289 42.74 0.9964 0.0362 40.77 0.9943 0.0408 39.39 0.9922 0.0447
HSDM-L 47.28 0.9984 0.0248 43.15 0.9965 0.0335 40.99 0.9946 0.0399 39.65 0.9924 0.0423
In this section, we explore synthetic datasets that are corrupted by mixed noise, real datasets, and remote datasets. We extensively assess our proposed HSDM and conduct experiments to analyze its efficacy thoroughly.
IV-A Overall Network Architecture
As illustrated in Figure 4, in the baseline configuration of HSDT, we employed 3 layers of HCSL, each comprising 4 Blocks for alternating forward and backward Continuous Scan. Additionally, before the modules, a residual connection and feature concatenation from the module output were utilized. Finally, skip connections were employed to add the initial input of each layer to the overall layer output, ensuring the stability of the denoising pattern.
IV-A1 Training Datasets
In training our proposed HSDM, we adopt a dataset configuration of [38], where 100 images are randomly sampled from the ICVL dataset [39]. These images were centrally cropped to achieve a spatial resolution of 512×512 with 31 spectral bands. Further adjustments are made to the spatial resolution of the images to 64×64 through operations including random cropping, rotation, and flipping, resulting in approximately 53,000 training samples.
IV-A2 Testing Datasets
To evaluate the model’s ability to suppress noise, testing datasets also encompass remote sensing datasets including the Washington DC dataset, real datasets[13] and Urban [41].
IV-A3 Comparison Methods and Evaluation Metrics
Our experiments encompass a comparative analysis between HSDM and a spectrum of methodologies, including traditional model-based approaches including LRTDTV [29] and LLRGTV [7], as well as deep learning-based methods such as GRNet [42], MACNET [31], T3SC [32], GRUNet[33], MAN [34], SST [35], SERT [36], and HSDT [37]. In particular, SST, SERT, and HSDT are Transformer-based methods, and others are hybrid-structured convolutional neural networks. To comprehensively evaluate the performance of these methodologies, we employ three distinct image quality evaluation metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM) [43], and Spectral Angle Mapper (SAM) [44]. Through this meticulous comparison, we aim to elucidate the superior denoising capabilities of HSDM in various real-world scenarios.
@@ IV-A3 Comparison Methods and Evaluation Metrics
HSDM-B 43.44 0.9967 0.0371 43.03 0.9964 0.0388 42.99 0.9964 0.0391 42.43 0.9958 0.0418 41.30 0.9950 0.0459
HSDM-L 43.84 0.9972 0.0347 43.25 0.9969 0.0366 43.23 0.9970 0.0354 42.55 0.9961 0.0396 41.62 0.9954 0.0431
Noisy LRTDTV LLRGTV MAC-Net T3SC
SST SERT HSDT HSDM-B GT
@@ IV-A5 Implementation Details
To align with the training configurations of mainstream approaches, for the training on Gaussian noise, we employed a mixture of Gaussian noises at varying levels. For complex noise, we introduced a random overlay of the first four types of noises, thereby enhancing the network’s adaptability and recovery capabilities. Additionally, HSDM utilizes the Adam[47] optimizer to minimize the mean squared error, following a training strategy similar to that outlined in [38], with minor modifications in the learning rate schedule. Specifically, the learning rate undergoes three stages: initially set to 3e-4, it is then reduced by a factor of 5 after every 20 epochs. Training is performed with a batch size of 8 on an RTX3090 for 100 epochs.
@@ IV-B Efficiency analysis
In this subsection, we analyze the performance of different methods regarding model parameter size and time consumption cost on the CAVE dataset with noise case 5. All test experiments are conducted on an RTX3090 GPU, and the results are presented in Table III.
TABLE III: Complexity comparisons on the CAVE dataset with dimensions of 512×512×31. Best in Blod, second best Underlined.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Noisy Noisy-zoom QRNN3D T3SC GRUNet SST SERT HSDT HSDM-B GT
@@ IV-C Experimental Results
To verify the efficacy of various comparative methods, we adopt a tailored approach. For model-based techniques, we carefully select test configurations that correspond to each model’s specific framework. In contrast, deep learning methodologies are evaluated against their publicly accessible pre-trained models, ensuring a fair comparison. The training procedure for all models utilizes the ICVL dataset to maintain consistency across evaluations.
Given the variable number of spectral bands across datasets, it is noteworthy that some comparative models are restricted to data comprising exactly 31 spectral bands. This limitation necessitates a selective application of these models based on dataset compatibility. To address the challenges posed by datasets with a broader spectral range, we implement a sliding-window strategy for the denoising task.
IV-C1 Gaussian Noise on Synthetic Datasets
In this section, we investigate the denoising performance of various methods under fixed Gaussian noise with different standard deviations ($\sigma$) ranging from 10 to 70, applied to the original HSI dataset. Our method is meticulously evaluated against baseline approaches on the ICVL dataset, with the comparative outcomes meticulously detailed in Table I. Under moderate and low levels of noise interference, our method exhibits performance comparable to SOTA methods. In particular, it demonstrates a denoising advantage when subjected to severe noise disturbances, particularly evident at the noise level of $\sigma=70$.
Figure 5 presents the denoising results derived from a variety of methods. A careful examination of these results reveals that methodologies such as LRTDTV, LLRGTV, and MAC-Net, despite their merits, display noticeable pseudo-artifacts in the presence of Gaussian noise. In stark contrast, our proposed HSDM method exemplifies remarkable performance in eliminating these Gaussian noise artifacts.
What sets HSDM apart is its ability to preserve the integrity of intricate details within the HSI dataset, a process deemed crucial for denoising. This not only enhances fidelity but also propels the accuracy to a notch higher compared to other available denoising methods. Manifestly, HSDM is adept at navigating the challenges posed by strong Gaussian noise and preserving intricate dataset details.
@@ IV-C2 Complex Noise on Synthetic Datasets
Furthermore, the results of visual comparison under mixed noise settings on the ICVL dataset are presented in Figure 6. It is noteworthy to observe a spectral shift phenomenon in denoising outcomes obtained with GRNet, wherein certain pixels transition from yellow to blue hues. Similarly, T3SC exhibits noticeable color distortion issues. In stark contrast, our proposed MambaHSID model maintains superior spectral consistency and visual fidelity, closely aligning with the GroundTruth across both spectral and spatial domains.
Figure 7: Comparative Band-by-Band Analysis of HSDM and Deep-Learning-Based Approaches on the WDC Dataset under Complex Noise Conditions.
Noisy
@@
LRTDTV LLRGTV MAC-Net T3SC QRNN3D
@@
GRUNet SST SERT HSDT HSDM-B
Figure 8: Visual comparison of denoising on the WDC dataset band 10 under mixture noise.
Noisy
@@
LRTDTV LLRGTV MAC-Net T3SC QRNN3D
@@
GRUNet SST SERT HSDT HSDM-B
Figure 9: Visual comparison of denoising on the WDC dataset band 76 under mixture noise.
Noisy
LRTDTV LLRGTV MAC-Net T3SC QRNN3D
GRUNet SST SERT HSDT HSDM-B
Figure 10: Visual comparison of denoising on the Urban dataset band 102 under mixture noise.
Noisy
LRTDTV LLRGTV MAC-Net T3SC QRNN3D
GRUNet SST SERT HSDT HSDM-B
Figure 11: Visual comparison of denoising on the Urban dataset band 138 under mixture noise.
Noisy
LRTDTV LLRGTV MAC-Net T3SC QRNN3D
GRUNet SST SERT HSDT HSDM-B
Figure 12: Visual comparison of denoising on the Urban dataset band 202 under mixture noise.
As shown in Figure 7, the comprehensive comparative analysis delineated through WDCMall under mixture noise shows that HSDM presents superior performance in a broad spectrum of bands. Figure 8 and Figure 9 elucidate the comparative efficacy of different bands, where ours consistently emerges as the most effective, indicating superior handling of variability and noise in the data.
@@ IV-C3 Noise on Real World
To validate the model’s robustness under real-world noise conditions, we conducted hyperspectral image denoising experiments on representative datasets Urban. Figures Figure 10, Figure 11, and Figure 12 illustrate the visual denoising performance across different bands of Urban datasets. Visual denoising performance is strikingly portrayed. In particular, our method demonstrates superior effectiveness, particularly in detail preservation and intensity transition regions. It meaningfully maintains the integrity of the details, subtly managing the noise and offering smoother transitions between diverse intensity levels. This superior demarcation of details during denoising and the exceptional performance at brightness transition zones underline the precision and robustness of our methodology.
Moreover, we extended our denoising experiments to include the CAVE dataset and the quantitative results are detailed in Table IV. HSDM showed remarkable superiority in real noise removal, as evidenced by its higher PSNR and SSIM values compared to other methods. Although LRTDTV and LLRGTV demonstrated a certain degree of alignment with the SAM metric changes on real datasets due to their priori global low-rank assumptions, their performance was relatively poorer on additional datasets. In contrast, HSDM consistently exhibited superior performance across a wider range of datasets, underscoring its versatility and adaptability to various HSI denoising scenarios.
@@ IV-D1 Effectiveness of scanning mechanism
To better explore the impact of scanning mechanisms combined with SSM on HSI denoising, we conducted experiments using both popular Sweep scanning and continuous scanning methods on different datasets. We performed both unidirectional and bidirectional experiments for each scanning method, and the results are summarized in Table IV, Alt representing whether to use the alternate mechanism. The findings provide evidence supporting the effectiveness of our bidirectional Continuous scanning approach.
@@
Baseline Alt Params ICVL CAVE WDC
@@
The results demonstrate that our bidirectional continuous scanning approach outperforms other scanning methods, indicating its effectiveness in capturing complex spatial-spectral relationships within the hyperspectral data. This finding underscores the importance of considering bidirectional scanning mechanisms when implementing state space models for HSI denoising.
@@ IV-D2 Effectiveness of modifications
IV-D3 Comparison with other vision mamba methods
In the chapter on dissolving experiment in our paper, we allocate a section to evaluate the performance of our proposition, HSDM, in comparison to the similar data-based VideoMamba and another similar denoising model, MambaIR. Unlike these two methodologies, HSDM employs a customized scanning architecture specifically designed for HSI and a pixel-smoothing scanning technique. This unique approach enables HSDM to achieve superior HSI denoising results at a more rapid pace shown in Table VII. Therefore, the proposed HSDM offers a compelling and efficient alternative in the HSI denoising domain, effectively dealing with challenges posed by other models while providing superior performance and speed in retrieval.
@@ IV-D3 Comparison with other vision mamba methods
hidden dim Params(M) Time(sec) PSNR(dB) SSIM SAM
MambaIR [20] 4.35 2.24 39.80 0.9900 0.0401
VideoMamba [19] 1.28 1.43 39.62 0.9843 0.0364
HSDM 0.68 0.77 41.30 0.9950 0.0459
IV-D4 Comparison with other architecture methods
To further elucidate the efficacy of our proposed HSDM architecture, we present a comprehensive comparison including train time(sec), inference time(sec), and GPU usage with GrUNet and Transformer-based methods on a 3090 GPU, the results are shown in Figure VI. The higher PSNR values of HSDM highlight its superior spatial-spectral utilization for HSI-denoising tasks. Enhanced computational efficiency and optimized utilization of GPU resources further position HSDM as an effective choice for comprehensive hyperspectral imaging endeavors. Thus, our HSDM validates the potency of the Mamba architecture within the HSI denoising domain.
TABLE VII: Ablation study of different architecture methods on CAVE dataset under mixture noise. SST and SERT are transformer-based methods.
Methods Params(M) Train Inference Usage PSNR
GRUNet [33] 41.44 0.27 0.63 9824 28.44
SST [12] 4.10 0.46 0.34 11348 34.89
SERT [13] 1.91 0.28 1.29 5104 35.86
HSDM 0.68 0.36 0.77 4644 41.30
@@ IV-E Parameter analysis
HSDM consistently maintains its GPU memory consumption and FPS rate. In contrast, the primary baseline Bidirectional HSDM experiences a decrease in FPS due to increased parameters. Notably, the impact of varying numbers of hidden channels on the model’s feature-capturing ability is significant. We analyze different numbers of hidden channels, presented in Table VIII. This decision strikes a better balance between the model’s efficiency and performance considerations. By increasing the number of hidden channels to 48, the model gains an enhanced ability to capture and represent complex features within the data. Thus, represents the HSDM-B version that we ultimately selected, while represent the HSDM-L version. This decision contributes to improved model performance without excessively increasing computational complexity, thus optimizing the trade-off between efficiency and effectiveness.
TABLE VIII: Ablation study of different spectral enhancement strategies on ICVL dataset under mixture noise.
Hidden dim Params(M) Runtime Usage PSNR(dB) SSIM SAM
24 0.24 0.58 2896 38.85 0.9901 0.0418
32 0.44 0.67 3766 39.71 0.9936 0.0350
48 0.68 0.79 4644 41.30 0.9950 0.0459
56 0.86 0.83 5282 41.50 0.9937 0.0374
64 1.08 0.90 5876 41.62 0.9954 0.0431
@@ V Conclusion
In this paper, we present the Hyperspectral Image Denoising Mamba-based network (HSDM), an innovative HSI denoising approach grounded in the selective state space model framework. HSDM employs a novel selective state-space model and a unique Hyperspectral Continuous Scanning Block (HCSB) for advanced denoising. HSDM effectively addresses the limitations of convolutional and attention mechanisms, leveraging an advanced 2D Selective Scan Module for spatial-spectral dependencies and maintaining linear complexity. The alternating continuous scanning mechanism improves efficiency and minimizes distortions. With the integration of residual learning and spectral attention, HSDM achieves exceptional noise reduction while preserving essential image features. Extensive experimental validations are conducted on diverse datasets, HSDM surpasses existing methods, highlighting its effectiveness and efficiency in tackling the challenges of HSI denoising.